• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 01:07:35 on August16,2001

    Entry number 70752

    keyword=magnets/ESR update

    ESR came back around midnight then down again after a few minutes.
    Recovered at about 0:20. Both spectrometer magenets levels were low at that
    time. Now the flows are fine, helium levels are recovering:
    Left arm Q2,Q3 back to normal (85%,95%), Q1, D are still 0%.
    Right arm Q2 is fine 94%, all others are low: Q3 is 30%, Q1 is 5%, D is 18%.